Slove each system of equations algebraically. x= ( y-3)^2 +2
y + x = 5

Answer by Nate(3500)

this is a "sideways" parabola and a linear equation
y + x = 5
y = -x + 5 get a variable by itself
------------------------------------------------------------------------------
x= ( y-3)^2 +2 original equation
x= (-x+2)^2 +2 fill in (-x+5) for y *see above*
x = x^2-4x+6 simplify
0 = x^2-5x+6 simplify
0 = (x-3)(x-2) simplify
x=3 these are possible x values in an ordered
x=2
------------------------------------------------------------------------------
y + x = 5 original equation
y + (3) = 5 plug in your x value
y = 2
(3,2)
y + x = 5 original equation
y + (2) = 5 5 plug in your other x value
y = 3
(2,3)